OpenStudy (jenny.andrade):

Determine the number of solutions you will have for the system 4y+x=16 and y=4-x

OpenStudy (anonymous):

a graph can be made from y=4-x

OpenStudy (mathstudent55):

The second equation is solved for y. Can you solve the first equation for y? What do you get?

OpenStudy (anonymous):

x=0. y=4

OpenStudy (mathstudent55):

No asked to solve the system. The only question is how many solutions there are.

OpenStudy (jenny.andrade):

is it no solutions, impossible to determine, one solution or indefinite solutions

OpenStudy (jdoe0001):

so what did you get for the 2nd equation when solved for "y"?

OpenStudy (mathstudent55):

@jenny.andrade Do you know how to solve the second equation for y?

OpenStudy (anonymous):

well its one set of solutions

OpenStudy (mathstudent55):

Here is the first equation: 4y + x = 16 We want to solve it for y. First, we subtract x from both sides: 4y = -x + 16 Now we divide both sides by 4: \(y = -\dfrac{1}{4}x + 4\) Now we bring in the first equation and we look at the system of equations with both equations solved for y. \(y = -\dfrac{1}{4}x + 4\) \(y=-x + 4\) We see that the equations are different. Therefore, we have two different lines. We also see that the slopes are different. That means the lines intersect in one single place. That means there is exactly one solution.

OpenStudy (jenny.andrade):

Nope. I don't know anything about algebra, so its going to be one solution?

OpenStudy (mathstudent55):

You could at least read what I wrote in response to your question.

OpenStudy (mathstudent55):

I just spent time trying to help you, but I know how valuable your time is, so if you'd like, just read the last line of my long answer above.

OpenStudy (anonymous):

@jenny.andrade its one sol.

OpenStudy (jenny.andrade):

thank you!!!!

OpenStudy (anonymous):

:)